OpenStudy (anonymous):

solve each system by graphing and tell whether it is consistent or inconsistent if the system is consistent tell whether it is dependent or independent x-y=6 5x-5y=30

OpenStudy (anonymous):

simplify the second equation

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

yes so divide the whole second equation by 5. what do you get?

OpenStudy (anonymous):

inconsistent, infinitly many solutions

OpenStudy (anonymous):

1x-1y=6

OpenStudy (anonymous):

you need to explain why @myko

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

yes so its the same as the other one right? :)

OpenStudy (anonymous):

let him think buy him self, :)

OpenStudy (anonymous):

Her

OpenStudy (anonymous):

can think too, no?

OpenStudy (anonymous):

yeah but she wanted help

OpenStudy (anonymous):

rebecca do you get it?

OpenStudy (anonymous):

i gave the answer

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

to solve equation of to unknowns you need at least 2 equations, but you only got one

OpenStudy (anonymous):

since they have_____________ __________ solutions.

OpenStudy (anonymous):

can you help me solve the other one and graph it please

OpenStudy (anonymous):

so you could give to one unknown any value and find the value for the other one, but in this way you can get infintly many pairs (x,y) that solve this equation, so kind of sad....

OpenStudy (anonymous):

they are the same equation so there is only one line. so therefore you need to only graph one line and there are infinetly many solutions.

OpenStudy (anonymous):

can you graph y=x-6?

OpenStudy (anonymous):

so how would i graph it i already got my equation so what to do i do

OpenStudy (anonymous):

the graphic solution of linear system of equations is two crossing lines. U only have one line that cross it self

OpenStudy (anonymous):

every point of the line is the answer to your system, :)

OpenStudy (anonymous):

1x-1y=6

OpenStudy (anonymous):

you get y by yourself and graph normally

OpenStudy (anonymous):

but what would the graph look like

OpenStudy (anonymous):



OpenStudy (anonymous):